# How to pick the best anomaly detector?

### v1

@@ A.2.2 HGB
HistGradientBoostingClassifier is a gradient boosted decision tree classifier implementation in scikit-learn [Pedregosa:2011sk], which histograms the classification features before training. It is currently state of the art for weakly supervised anomaly detection [Finke:2023ltw, Freytsis:2023cjr] with high-level feature. Following Ref. [Finke:2023ltw], we perform 50 independent trainings with different 50-50 training-validation splits, which we ensemble by taking the mean over the predictions. Our default hyperparameters are largely equivalent to the default of the scikit-learn package and can be found in Tab. 2. For hyperparameters optimization, we sample hyperparameters from the distributions also shown in the table.
@@ A.2.3 AdaBoost
AdaBoostingClassifier is an adaptive boosted decision-tree classifier implementation in scikit-learn, which was investigated as an alternative to HGB in [Finke:2023ltw] but did not achieve the performance or robustness of HGB. We build our implementation in analogy to [Finke:2023ltw] and perform an ensembling of 10 independently trained classifiers using random 50-50 splits of the training dataset. AdaBoost does not make use of a validation set. However, as training is deterministic, the random data split results in variations between different trainings, which can be used to obtain a better performance through ensembling. Our default hyperparameters reflect the default used for AdaBoost in [Finke:2023ltw] and can be found in Tab. 3. The distributions used to randomly sample hyperparameters for optimization can be found there, as well.
@@ Appendix C Details of the architecture selection
The first step of our architecture selection is to train our architecture options – NN, HGB and AdaBoost, on 50% of our normal training and validation set. The anomaly detection performance characterized by the max SIC of these runs is shown in the middle panel of Fig. 7. On the other half of the data, we then calculate the ARGOS metric, the result of which is shown in the left panel of Fig. 7. For both of these panels we show the median and 16 to 84 percent quantiles of ten runs as lines and error bands respectively. These ten runs contain not only ten independent trainings but also ten different hyperparameter configurations per model as we use the results from Sec. IV.3. For each of these sets of ten runs, we then pick the best-performing architecture between NN, HGB and AdaBoost based on the obtained max ARGOS values and plot the resulting set of ten again as median and quantile-error band. We perform this 50-50 split of the data set in order to have the metric values used for architecture selection be calculated on a held-out evaluation set. As all classifiers use the validation set in one form or another, this is not the case when directly selecting on the validation set.

### v2

@@ A.2.2 HGB
HistGradientBoostingClassifier is a gradient boosted decision tree classifier implementation in scikit-learn [Pedregosa:2011sk], which histograms the classification features before training. It is currently state of the art for weakly supervised anomaly detection [13, 14] with high-level feature. Following Ref. [13], we perform 50 independent trainings with different 50-50 training-validation splits, which we ensemble by taking the mean over the predictions. Our default hyperparameters are largely equivalent to the default of the scikit-learn package and can be found in Tab. 2. For hyperparameters optimization, we sample hyperparameters from the distributions also shown in the table.
@@ A.2.3 AdaBoost
AdaBoostingClassifier is an adaptive boosted decision-tree classifier implementation in scikit-learn, which was investigated as an alternative to HGB in [13] but did not achieve the performance or robustness of HGB. We build our implementation in analogy to [13] and perform an ensembling of 10 independently trained classifiers using random 50-50 splits of the training dataset. AdaBoost does not make use of a validation set. However, as training is deterministic, the random data split results in variations between different trainings, which can be used to obtain a better performance through ensembling. Our default hyperparameters reflect the default used for AdaBoost in [13] and can be found in Tab. 3. The distributions used to randomly sample hyperparameters for optimization can be found there, as well.
@@ Appendix C Details of the architecture selection
Figure 8: Background sculpting for IAD (left), CATHODE (middle) and CWoLa Hunting (right) when selecting epochs based on validation BCE loss, max ARGOS or selecting the last epoch characterised by the $\chi^{2}$ metric from [20] (top) and example spectra obtained when selecting $1\%$ most anomalous events (bottom).
The first step of our architecture selection is to train our architecture options – NN, HGB and AdaBoost – on 50% of our normal training and validation set. The anomaly detection performance characterized by the max SIC of these runs is shown in the middle panel of Fig. 7. On the other half of the data, we then calculate the ARGOS metric, the result of which is shown in the left panel of Fig. 7. For both of these panels we show the median and 16 to 84 percent quantiles of ten runs as lines and error bands respectively. These ten runs contain not only ten independent trainings but also ten different hyperparameter configurations per model as we use the results from Sec. IV.3. For each of these sets of ten runs, we then pick the best-performing architecture between NN, HGB and AdaBoost based on the obtained max ARGOS values and plot the resulting set of ten again as median and quantile-error band. We perform this 50-50 split of the data set in order to have the metric values used for architecture selection be calculated on a held-out evaluation set. As all classifiers use the validation set in one form or another, this is not the case when directly selecting on the validation set.
@@
Appendix D False discoveries and background sculpting
When optimizing on data, it is important to understand the effect that this has on the statistics of an analysis. In our case, we seek to understand whether an increased rate of false discoveries may follow, which could stem from one of two issues:
  1. 1.
    Selecting the best classifier on data could lead to “p-hacking”: The selected classifier may prioritize purely statistical fluctuations in the training data, leading to a look elsewhere effect [Hein:2025ysv] and miscalibrated significances.
  2. 2.
    The selected classifier may prioritize regions of the data with the most systematic background mismodeling, which could lead to false excesses in both fit-based bump hunts (as performed in Refs. [ATLAS:2020iwa, 8, 1, 15]) and in analyses with direct background estimation [Das:2024eie].
The first issue can be completely eliminated by evaluating on an independent test set. This will be the case for the remainder of this appendix.
Regarding the second issue, the classifier most able to distinguish between data and background is fundamentally the better anomaly detector. If this leads to false positives due to background mismodeling, this is not a fault of the anomaly detector per se.
Nevertheless, we will investigate this issue in more detail, to see what the effect of improved model selection has on it.
Shown in Fig. 8 is the $\chi^{2}$ histogram difference for the background $m_{JJ}$ distribution before and after a cut on the anomaly score, as a function of background selection efficiency. This metric was previously introduced in Ref. [20] and is a way to quantify the effect of background sculpting due to a cut on the anomaly score. We compare the sculpting for an anomaly detector selected using the BCE loss (as used in Ref. [8]), the ARGOS metric and using the last epoch. We see that for the IAD and CATHODE the sculpting does not change significantly based on the different model selection methods.
For CWoLa Hunting, the amount of sculpting is already more severe compared with IAD or CATHODE when model selecting with BCE loss or using the last epoch. This is further exacerbated when the selection is performed with ARGOS. This is entirely expected as the background in the SR is not perfectly modeled by the SB data, and with the ARGOS metric we are picking this up more effectively.
If significances are obtained through direct background estimation the magnitude of the systematic shift between BT and SR needs to be estimated in any case for the given classifier setup [Das:2024eie]. Using a better classifier based on ARGOS optimization, the systematic shift might increase. However, estimating the systematic shift properly, an increased rate of false discoveries will be avoided.